4. Church Inc. is presently enjoying relatively high growth because of a surge in the demand for its new product. Management expects earnings and dividends to grow at a rate of 25% for the next 4 years, after which competition will probably reduce the growth rate in earnings and dividends to zero, i.e., g = 0. The company’s last dividend, D0, was $1.25, its beta is 1.20, the market risk premium is 5.50%, and the risk-free rate is 3.00%. What is the current price of the common stock?

a. $26.77
b. $27.89
c. $29.05
d. $30.21
e. $31.42
1
Expert's answer
2012-01-27T08:27:12-0500
a. $26.77
Pa = d0*(1+g/100)*100/(k - g) = 1,25*(1+25/100)*100/6 = 26.77